OpenStudy (anonymous):

Verify the identity tan x+pi/2 =-cot x please?

OpenStudy (anonymous):

pleaaaaase :)

OpenStudy (ddcamp):

This would be a good place to start: \[\tan(x+\frac{ \pi }{ 2 }) = \frac{ \sin(x+\frac{ \pi }{ 2 }) }{ \cos(x+\frac{ \pi }{ 2 }) }\]

OpenStudy (anonymous):

how about using the sum identity?

OpenStudy (anonymous):

sorry, not necessary

OpenStudy (anonymous):

so from what he said don't you have to relate that to -cotx?

OpenStudy (anonymous):

DDCamp has it right.

OpenStudy (anonymous):

\[\cos \left( x+\frac{\pi}{2} \right)=-\sin x\] \[\sin \left( x+\frac{\pi}{2} \right)=\cos x\]

OpenStudy (ddcamp):

\[\cot(x)=\frac{ \cos(x) }{ \sin(x) }\]

OpenStudy (anonymous):

and then isn't it a Co-Function Identities like cot= tan(x+pi/2)

OpenStudy (anonymous):

oh never mind! Thank you @DDCamp you rock and saved my butt :)

OpenStudy (anonymous):

and @pgpilot326 you rock too sorry :D

OpenStudy (anonymous):

\[\tan \left( x+\frac{ \pi }{ 2 } \right)=\frac{ \sin \left( x+\frac{ \pi }{ 2 } \right) }{ \cos \left( x+\frac{ \pi }{ 2 } \right)}=\frac{ \cos x }{ -\sin x }=-\cot x\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@pgpilot326 got it thank you so much :) I'll give the last answer best answer :)

OpenStudy (anonymous):

no worries!